# SALSA-RL: Stability Analysis in the Latent Space of Actions for Reinforcement Learning

### v1

@@ A.2 Policy Gradient
Term 1: $\partial\,\text{Dec}(z)/\partial z$. For Decoder that is fixed/pre-trained, this Jacobian is a constant.
Term 2: $\nabla_{\theta}\,h_{\theta}(s_{t},a_{t})$. From Eq. (17),
@@
Appendix B Algorithms for Transient Growth and Floquet Analysis
@@
Transient Growth. For transient growth analysis in Section 3.5, we describe the calculation of the Kreiss constant. Regarding the choice of $z$ sampling, which produces the results in Figure 5, the upper bound is set to 1.5 for the first pendulum case and 8 for the second.
@@
[…]
@@
Floquet Analysis. For Floquet Analysis in Section 3.5, the algorithm is summarized below:
@@ Appendix E Extended Stability Analysis and Interpretability on Pendulum Problem
In the main article, the analysis of pendulum control highlighted regions of stability along with the control actions over time. in Figure 6, we examine extreme scenarios involving the absence of actions as potential control failures, aiming to showcase extended stability analysis during these failure processes.
Figure 6: A custom control scenario where the system remains action-free during the first 30 and last 50 timesteps. The first column shows the evolution of action, state, and corresponding eigenvalues, while the second and third columns present an extended stability analysis through contour visualizations.
@@
    Region of Instability (T: 0–30): Initial instability before any actions are applied.
@@
    Region of Recovery (T: 30–150): Control actions are applied to recover the system from instability to stability.
@@
    Region of Control Failure (T: 150–200): The control is removed, and the system transitions back from stability to instability.
@@
From the eigenvalue evolution plot (first column, third row), regions of instability ($\lambda_{1}>1$) are evident in time Regions 1 and 3, as well as at the initial stage of Region 2. In Region 2, the policy successfully recovers the system from instability to stability, aligning with the above-defined phases and the pendulum control objective. In the second and third columns, contour plots further validate this observation. Contour regions of instability or oscillation near the trajectory are visibly highlighted with yellow or red colors, reinforcing our stability analysis and definition. Overall, this analysis significantly enhances our understanding of control failures, showcasing the interpretability and robustness of the proposed framework.
Appendix F Stability Analysis for the BipedalWalker Control
@@
The Bipedal Walker problem involves numerous states and four actions controlling the legs and joints, aiming to achieve stable, energy-efficient forward locomotion across uneven terrain without falling. Similar to other environments analyzed in the main article, we present the stability analysis for two different cases, as shown in Figure 7. Due to the large number of states, we illustrate only four representative states associated with the walker’s hull.
@@
Figure 7: Stability analysis of the Bipedal Walker control for two cases (left two columns and right two columns). Representative states associated with the walker’s hull are shown. Both cases exhibit similar control patterns and stability. The observed states (red lines) indicate that the hull maintains a positive angle, while angular velocity and vertical velocity (second row) oscillate around 0. The horizontal velocity (third row) remains positive, enabling forward movement. These observed states closely align with the stability contours. The real part of the eigenvalues stays mostly below 1, indicating stability, while the first three plots illustrate how the control policy stabilizes the states within potential instability regions (white-line regions). The imaginary part, with magnitudes greater than 1, reflects oscillatory dynamics in the overall states.
In both cases, similar stability patterns emerge over time. The hull primarily maintains stability through oscillations, meaning the real part of the eigenvalues generally remains below 1, indicating stability, while the imaginary part remains consistently nonzero, reflecting oscillatory dynamics. This corresponds to the walker’s steady rightward movement, with occasional hull oscillations (moving up and down) as the legs alternate between stance and swing phases.
Notably, in the second row (angular velocity contour), regions around the state trajectory indicate high values (instability). This corresponds to the walker, where high angular velocity (rotational speed) of the hull signifies instability. In the third row (horizontal velocity contour), negative hull velocity represents instability, as the walker should move to the right rather than the left, aligning with the environment’s objective. In contrast, in the first row (angle of the hull), variations within a certain range are acceptable for maintaining balance. While in the fourth row (vertical velocity contour), values remain small and stable, reflecting consistent vertical motion. Other states related to the legs, which exhibit constant movement and consistently show regions of instability, are excluded from the analysis as they are not of primary interest.
@@ Appendix G Phase-Space Visualizations of SALSA-RL
For the pendulum problem, where the state space is summarized by two components–the angle $\theta$ and angular velocity $\dot{\theta}$, rather than $\theta$, $\cos\theta$, and $\sin\theta$–we analyze the phase-space behavior of the learned policies. By interacting the policy with the environment, we visualize the structure of actions assigned to different states. Figure 8 compares the baseline SAC [35] with the proposed SALSA-RL, both sharing similar architectures and training strategies. On the right, the SAC policy results in scattered action assignments across phase space, reflecting its probabilistic nature. In contrast, on the left, SALSA-RL exhibits larger, more coherent action regions, indicating more consistent action choices for given states. This suggests that SALSA-RL encodes a structured latent representation, leading to more predictable decision-making in phase space.
@@ Appendix H Ablation Study on Stability Analysis Across Different Hidden Dimensions
In Figure 9, we extensively show the effect of different hidden dimensions contributing to the stability analysis. In the upper row with lower $h_{d}$, the contour patterns are highly similar within the real component plots and within the imaginary component plots. This consistency aligns with the main article and is particularly evident in the initial stage, where the control policy seeks stability while crossing regions of instability.
@@
In the lower row with higher $h_{d}$, the evolution of the real part exhibits different values. However, the overall pattern remains consistent across different $h_{d}$ configurations, as the trajectory initially passes through yellow-colored regions. Notably, at $h_{d}=8$, the system exhibits uniform stability across the entire domain, whereas at $h_{d}=16$, regions of instability emerge. It is worth noting that in the higher $h_{d}$ cases, even after the pendulum stabilizes (beyond 50 timesteps), the imaginary component reveals regions of high values, indicating oscillatory behavior. However, these regions remain distant from the actual state of the pendulum, and the increased values are likely due to the higher-dimensional neural network introducing greater nonlinearity.
@@
Figure 9: Stability contour plots for different hidden dimensions: (top-left) $h_{d}=3$, (top-right) $h_{d}=4$, (bottom-left) $h_{d}=8$, and (bottom-right) $h_{d}=16$. Despite some discrepancies in numerical values—such as the real part in $h_{d}=8$ showing uniform stability—similar patterns and overall behaviors hold across different configurations. This consistency reinforces the robustness of our stability analysis, demonstrating that key trends persist regardless of the hidden dimension size.

### v3

@@ A.2 Policy Gradient
Term 1: $\partial\,\text{Dec}(z)/\partial z$. For Decoder that is fixed/pre-trained, this Jacobian is a constant. Term 2: $\nabla_{\theta}\,h_{\theta}(s_{t},a_{t})$. From Eq. (17),
@@
Appendix B Algorithms for Local Stability Analysis, Transient Growth, and Floquet Analysis
@@
Algorithm 2 Local Stability Evaluation
Input: Observed $n$ states $s_{t}^{(i)}$, where i=1,2,,ni=1,2,...,n
for $i=1$ to $n$ do
  for $t=0$ to $T$ do
   for $s$ in $s_{\text{range}}^{(i)}$ do
    $\hat{\mathbf{s}}[i]\leftarrow s$ // Replace with neighboring values
[…]
    $\lambda_{i}\leftarrow\text{eigvals}(\mathbf{A}_{t})$ // Eigenvalue calculation
    $\rho(\mathbf{A}_{t})\leftarrow\max_{i}\left|\lambda_{i}\right|$ // Spectral radius
    $\mathrm{Im}_{\max}\leftarrow\max\left(\mathrm{Im}(\lambda_{i})\right)$ // Maximum imaginary magnitude
   end for
  end for
end for
@@
   $\text{normality\_diff}\leftarrow\|\mathbf{A}_{t}\cdot\mathbf{A}_{t}^{\top}-\mathbf{A}_{t}^{\top}\cdot\mathbf{A}_{t}\|$
   if $\text{normality\_diff}>\epsilon$ then
@@ Appendix E Extended Stability Analysis and Interpretability on Pendulum Problem
In the main article, the analysis of pendulum control highlighted regions of local stability along with the control actions over time. In Figure 6, we examine extreme scenarios involving the absence of control actions as potential control failures, aiming to showcase extended stability analysis during these failure processes.
Figure 6: A custom control scenario where the system remains action-free during the first 30 and last 50 timesteps. The first column shows the evolution of action, state, and corresponding spectral radius and imaginary part of the eigenvalues, while the second and third columns present an extended stability analysis through contour visualizations.
@@
    Region of Instability (T: 0–30): Initial local instability before any actions are applied.
@@
    Region of Recovery (T: 30–150): Control actions are applied to recover the system from local instability to stability.
@@
    Region of Control Failure (T: 150–200): The control is removed, and the system transitions back from local stability to instability.
@@
From the spectral radius and eigenvalue plot (first column, third row), regions of local instability ($\rho_{1}>1$) are evident in time Regions 1 and 3, as well as at the initial stage of Region 2. In Region 2, the policy successfully recovers the system from local instability to stability, aligning with the above-defined phases and the pendulum control objective. For the imaginary part of the eigenvalue evolution, pronounced oscillatory behavior is observed in Regions 1 and 3, which aligns with the corresponding periods of local instability identified by the spectral radius analysis.
In the second and third columns of the figure, contour plots further validate this observation. Regions of local instability or oscillation near the trajectory are visibly highlighted in yellow and red. These appear on both the left and right sides of the contour plots, corresponding to time regions 1 and 3, and reinforce our earlier local stability analysis and definition. Overall, this analysis significantly enhances our understanding of control failures, showcasing the interpretability and robustness of the proposed framework.
Appendix F Local Stability Analysis for the BipedalWalker Control
@@
The Bipedal Walker problem involves numerous states and four actions controlling the legs and joints, aiming to achieve stable, energy-efficient forward locomotion across uneven terrain without falling. Similar to other environments analyzed in the main article, we present the local stability analysis for two different cases, as shown in Figure 7. Due to the large number of states, we illustrate only four representative states associated with the walker’s hull.
@@
Figure 7: Local stability analysis of the Bipedal Walker control for two cases (left two columns and right two columns). Representative states associated with the walker’s hull are shown. Both cases exhibit similar control patterns and local stability. The observed states (red lines) indicate that the hull maintains a positive angle, while angular velocity and vertical velocity (second row) oscillate around 0. The horizontal velocity (third row) remains positive, enabling forward movement. The large variations in angle and vertical velocity align closely with the local instability observed in the spectral radius contour, which is primarily driven by strong oscillations, as indicated by the imaginary eigenvalue contour.
In both cases, the hull primarily exhibits local instability throughout most of the episode, with the spectral radius remaining significantly greater than 1. These high values are driven largely by the magnitude of the imaginary components of the eigenvalues, even though the real parts remain mostly below 1. From the walker’s perspective, this instability, caused by persistent oscillations, corresponds to the agent’s continuous rightward movement. The oscillatory behavior of the hull (i.e., vertical up-and-down motion) emerges naturally as the legs alternate between stance and swing phases. This is also evident in the angle and angular velocity states (shown in the first and second rows), where significant variations in the hull’s orientation are observed.
Additionally, during movement, the vertical velocity of the hull (shown in the last row of Figure 7) is minimized and remains close to zero. This behavior is also reflected in the imaginary eigenvalue contour, where the red trajectory resides in a region of low imaginary components (depicted in blue). However, oscillations during movement cannot be entirely eliminated, resulting in non-zero imaginary eigenvalues and, consequently, a system characterized by $\rho>1$ and continuous use of actions for controlling the system. This example reflects how the spectral radius must be interpreted in the context of the state-action behavior of a system being controlled.
@@ Appendix G Phase-Space Visualizations of SALSA-RL
For the pendulum problem, where the state space is summarized by two components–the angle $\theta$ and angular velocity $\dot{\theta}$, rather than $\theta$, $\cos\theta$, and $\sin\theta$–we analyze the phase-space behavior of the learned policies. By interacting the policy with the environment, we visualize the structure of actions assigned to different states.
Figure 8 compares the baseline SAC [42] with the proposed SALSA-RL, both sharing similar architectures and training strategies. On the right, the SAC policy results in scattered action assignments across phase space, reflecting its probabilistic nature. In contrast, on the left, SALSA-RL exhibits larger, more coherent action regions, indicating more consistent action choices for given states. This suggests that SALSA-RL encodes a structured latent representation, leading to more predictable decision-making in phase space. These slight discrepancies arise from the encoder-decoder structure in SALSA-RL, which transforms raw actions into a learned latent space before applying them, thus altering the effective action dynamics. Notably, this transformation may also enhance the stability of the pretrained agent, though a full analysis of this effect is left for future work. Like [53], our approach may also be used to perform principled data augmentation for improved behavior of RL agents in terms of regularity.
@@ Appendix H Ablation Study on Stability Analysis Across Different Hidden Dimensions
In Figure 9, we extensively show the effect of different hidden dimensions contributing to the stability analysis. Overall, the stability patterns remain consistent across different $h_{d}$ configurations: the trajectory initially passes through high-instability regions (yellow) before stabilizing over time.
@@
In the first row with low $h_{d}$, the contour patterns are highly similar within the spectral radius and within the imaginary component plots. This consistency aligns with the main article and is particularly evident in the initial stage, where the control policy seeks local stability while crossing regions of instability. In the second row with high $h_{d}$, the spectral radius contour reveals more symmetric and well-defined regions of stability, in contrast to the upper row, where local stability is only observed on the lower-value side and the upper region remains unbounded or weakly constrained. This suggests that higher latent dimensions enable more accurate local stability characterization, particularly evident in the angular velocity plot.
@@
It is worth noting that in the higher $h_{d}$ cases, even after the pendulum stabilizes (beyond 50 timesteps), the imaginary component reveals regions of high values, indicating oscillatory behavior. However, these regions remain distant from the actual state of the pendulum, and the increased values are likely due to the higher-dimensional neural network introducing greater nonlinearity.
@@
Figure 9: Local stability contour plots for different hidden dimensions: (top-left) $h_{d}=3$, (top-right) $h_{d}=4$, (bottom-left) $h_{d}=8$, and (bottom-right) $h_{d}=16$. Despite some discrepancies in numerical values, such as the real part in $h_{d}=8$ showing uniform stability, similar patterns and overall behaviors hold across different configurations. This consistency reinforces the robustness of our local stability analysis, demonstrating that key trends persist regardless of the hidden dimension size.
Appendix I Stability Analysis on High-Dimensional Humanoid Environments
To further assess robustness, we extend SALSA-RL to high-dimensional continuous control, specifically the Humanoid task in Isaac Gym [36]. This environment features an action space of 21 dimensions and an observation space of 108 dimensions.
For the autoencoder, we train models with latent dimensions $h_{d}$ in higher values of $\{32,64,128\}$, since lower values significantly degrade reconstruction performance. Both the encoder and decoder are implemented as three-layer neural networks with 256 hidden units per layer.
For reinforcement learning, we adopt PPO with a policy network consisting of three layers of 400, 200, and 100 hidden units, respectively. Training is performed for 3000 epochs. Standard PPO achieves an average reward of approximately 10447 for 1000 timesteps. In comparison, SALSA-RL achieves 8999 for $h_{d}=32$, 9390 for $h_{d}=64$, and 9170 for $h_{d}=128$.
Representatively, we show the spectral radius and imaginary eigenvalue trajectories over time for the case $h_{d}=32$, as illustrated in Figure 10. The spectral radius fluctuates around the unit boundary, with frequent crossings above and below $\rho(\mathbf{A}_{t})=1$. This indicates that the learned policy operates at the edge of stability, exhibiting persistent transitions between stable and unstable regimes in the latent action dynamics. While transient excursions into instability are observed, the spectral radius tends to remain close to 1, suggesting that the overall behavior of the policy is near-critical but not dominated by instability.
In the extended analysis, we visualize the stability contours over time for the first dimension of the observation space (e.g., the first state variable), as shown in the second row of Figure 10. These plots indicate that, beyond growth rates captured by the spectral radius, imaginary components make notable contributions, reflecting oscillatory modes that influence the observed stability transitions.
Overall, the humanoid results show that SALSA-RL achieves comparable performance to PPO while additionally enabling post-hoc stability analysis in high-dimensional settings. This demonstrates the scalability and robustness of the framework, highlighting its potential for analyzing complex control policies beyond low-dimensional benchmark tasks.
Figure 10: Stability analysis for Humanoid with $h_{d}=32$. Top row: evolution of actions, states, spectral radius, and imaginary parts of eigenvalues. Bottom row: extended analysis via stability contour visualizations.
Appendix J SALSA-RL Complexity Analysis
For a standard two-layer actor with state dimension $n$, action dimension $m$, and hidden width $H$, the computational complexity is
$\mathcal{C}_{\text{Actor}}=O(nH+H^{2}+Hm).$ (23)
For SALSA-RL, the complexity includes the latent dynamics matrix $A_{t}$, latent updates, and the encoder/decoder modules:
$\mathcal{C}_{\text{SALSA-RL{}}}=O(nH+H^{2}+Hh_{d}^{2}+h_{d}^{2}+2(mc+c^{2}+ch_{d})),$ (24)
where $h_{d}$ is the latent dimension and $c$ is the hidden width of the two-layer encoder/decoder.
Since $h_{d}\ll H\approx c$ in all our experiments (e.g., $h_{d}\leq 16$ while $H,c\approx 200$), the additional terms are negligible relative to the dominant $H^{2}$ term of the baseline. Hence, the overall complexity of SALSA-RL simplifies to
$\mathcal{C}_{\text{SALSA-RL{}}}\approx O(H^{2}+Hh_{d}^{2}).$ (25)